[Hallc_running] [New Logentry] Beam status for tonight

hszumila at jlab.org hszumila at jlab.org
Thu Feb 8 18:45:01 EST 2018

Logentry Text:
Just spoke with PD. The machine is currently very unstable and unable to deliver much current to Halls A & C. Looks like the max we can expect tonight in C will be around 26 uA or so. We will have to see. Unfortunately, due to the instability, the preliminary plan is to take the accelerator down tomorrow morning at the start of day shift to do a complete re-tooling of the machine. This is expected to take 3 days (Friday-Sunday). Details will be discussed and decided tomorrow morning at the usual MCC meeting. In the meantime, all shifts should continue as usual until we have more information at the meeting tomorrow morning at which time, we will know about the fate of the later shifts into the weekend. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3529376
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180208/91ec8c3e/attachment.html>

More information about the Hallc_running mailing list